### v1

@@ 5 Error Estimators for Non-symmetric Problems
A rigorous study of smoother-type a posteriori error estimates for non-symmetric problems is beyond the scope of the current paper. Nevertheless, error estimators such as $\|S^{a}_{h/2}r\|$ is still applicable to arbitrary linear PDEs.
@@ 5.1 Convection-diffusion equation
(5.1) $-\nabla\cdot(\alpha\nabla u+\beta u)$ $=f\quad\text{ in }\partial\Omega,$
@@
The polynomial degree $p$ ranges from 1 to 7. The initial mesh size $h_{0}$ is set as $h_{0}=2^{-7}$ for $p\leq 2$ and $h_{0}=2^{-5}$ for $p>2$.
The experimental results in Figure 12 demonstrate that the smoother-type estimators remain to be effective for the highly non-symmetric problem (5.1). As shown in Figure 11, the corresponding AFEM is able to capture both the interface and boundary singularity. It is also observed from Figure 12 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ remain to be mild as the polynomial degree $p$ grows.
Figure 11: Adaptive meshes with approximately $2.2\times 10^{4}$ triangles produced by $\mathcal{P}_{1}$- (left) and $\mathcal{P}_{5}$-AFEM (right) driven by the Gauss–Seidel smoother-type error estimator $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$.
Figure 12: Convergence of $\mathcal{P}_{p}$-AFEM for the convection-diffusion problem driven by Jacobi (Left) and Gauss–Seidel (Right) smoother estimators with ratio estimator/error shown in the parenthesis.
@@ 5.2 Helmholtz Equation
(5.2a) $-\Delta u-k^{2}u$ $=f\quad\text{ in }\Omega,$
(5.2b) $\partial_{n}u-\texttt{i}ku$ $=g\quad\text{ on }\partial\Omega_{1},$
@@
Here $(\bullet,\bullet)$ and $(\bullet,\bullet)_{\partial\Omega_{1}}$ are complex-valued inner products. A natural norm in the error analysis of (5.2) is $\|v\|_{1,k}=\sqrt{k^{2}\|v\|_{L^{2}(\Omega)}^{2}+|v|_{H^{1}(\Omega)}^{2}}$ (cf. [55, 33]). Similarly to section 5.1, we use the Jacobi $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}$ and Gauss–Seidel $\|S^{m}_{\mathcal{P}_{p}}r\|_{1,k}$ a posteriori error estimates of AFEMs for (5.2).
@@
is a cutoff function. The polynomial degree $p$ ranges from 1 to 7 and the initial mesh size $h_{0}$ is $h_{0}\approx 1/60$ for $p=1$ and $h_{0}\approx 1/30$ for $p>1$. The experimental results in Figure 14 demonstrate that the smoother-type estimators remain to be effective for the Helmholtz equation with large wave number $k=40\pi$. As shown in Figure 13, AFEMs based on smoother-type a posteriori error estimates correctly capture the oscillation and corner singularity of (5.2). It is also observed from Figure 14 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ are similar and uniform as the polynomial degree $p$ grows. Considering the computational cost, $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ is more efficient than its Gauss–Seidel counterpart.
Figure 13: Adaptive meshes (approximately $10^{5}$ triangles) by $\mathcal{P}_{1}$-AFEM (left) and $\mathcal{P}_{4}$-AFEM (right) driven by the pointwise Jacobi smoother error estimator $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$.
Figure 14: Convergence of $\mathcal{P}_{p}$-AFEM driven by Jacobi (Left) and Gauss–Seidel error estimators (Right) for the Helmholtz Equation with $k=40\pi$ and estimator/error shown in the parenthesis.
@@ 6 Concluding Remarks
This paper bridges a gap between a posteriori error estimates and iterative linear solvers. Leveraging smoothers of available preconditioners and linear solvers, we have derived a posteriori error estimates for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. These estimators are computationally efficient, parameter-robust, polynomial-degree-robust and are validated through extensive numerical experiments.
@@ Acknowledgments
Y. L. would like to thank Prof. Shihua Gong for helpful discussion about iterative solvers for the Helmholtz equation.

### v2

@@ 5 Error Estimators for Non-symmetric Problems
A rigorous study of smoother-type a posteriori error estimates for non-symmetric problems is beyond the scope of the current paper. Nevertheless, error estimators such as $\|S^{a}_{h/2}r\|$ are still applicable to arbitrary linear PDEs.
@@ 5.1 Convection-diffusion equation
$-\nabla\cdot(\alpha\nabla u+\beta u)$ $=f\quad\text{ in }\partial\Omega,$ (5.1)
@@
The polynomial degree $p$ ranges from 1 to 7. The initial mesh size $h_{0}$ is set as $h_{0}=2^{-7}$ for $p\leq 2$ and $h_{0}=2^{-5}$ for $p\geq 3$, as low order AFEMs starting from a very coarse initial grid fail to converge.
The experimental results in Figure 5.2 demonstrate that the smoother-type estimators remain effective for the highly non-symmetric problem (5.1). As shown in Figure 5.1, the corresponding AFEM is able to capture both the interface and boundary singularity. It is also observed from Figure 5.2 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ remain mild as the polynomial degree $p$ grows.
Figure 5.1: Adaptive meshes with approximately $2.2\times 10^{4}$ triangles produced by $\mathcal{P}_{1}$- (left) and $\mathcal{P}_{5}$-AFEM (right) driven by the Gauss–Seidel smoother-type error estimator $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$.
Figure 5.2: Convergence of $\mathcal{P}_{p}$-AFEM for the convection-diffusion problem driven by Jacobi (Left) and Gauss–Seidel (Right) smoother estimators with ratio estimator/error shown in the parenthesis.
Figure 5.3: Effectivity ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u_{h}-u_{h/2}|_{H^{1}(\Omega)}$ (top) and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u_{h}-u_{h/2}|_{H^{1}(\Omega)}$ (bottom) for the convection-diffusion equation.
@@ 5.2 Helmholtz Equation
$-\Delta u-k^{2}u$ $=f\quad\text{ in }\Omega,$ (5.2a)
$\partial_{n}u-\texttt{i}ku$ $=g\quad\text{ on }\partial\Omega_{1},$ (5.2b)
@@
Here $(\bullet,\bullet)$ and $(\bullet,\bullet)_{\partial\Omega_{1}}$ are complex-valued $L^{2}$ inner products. A natural norm in the error analysis of (5.2) is $\|v\|_{1,k}=\sqrt{k^{2}\|v\|_{L^{2}(\Omega)}^{2}+|v|_{H^{1}(\Omega)}^{2}}$ (cf. [64, 41]). Similarly to Section 5.1, we use the Jacobi $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}$ and Gauss–Seidel $\|S^{m}_{\mathcal{P}_{p}}r\|_{1,k}$ a posteriori error estimates of AFEMs for (5.2).
@@
is a cutoff function. The polynomial degree $p$ ranges from 1 to 7 and the initial mesh size $h_{0}$ is $h_{0}\approx 1/60$ for $p=1$ and $h_{0}\approx 1/30$ for $p\geq 2$. The experimental results in Figure 5.5 demonstrate that the smoother-type estimators remain effective for the Helmholtz equation with large wave number $k=40\pi$. As shown in Figure 5.4, AFEMs based on smoother-type a posteriori error estimates correctly capture the oscillation and corner singularity of (5.2). It is also observed from Figure 5.5 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ are similar and uniform as the polynomial degree $p$ grows. Considering the computational cost, $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ is more efficient than its Gauss–Seidel counterpart.
Figure 5.4: Adaptive meshes (approximately $10^{5}$ triangles) by $\mathcal{P}_{1}$-AFEM (left) and $\mathcal{P}_{4}$-AFEM (right) driven by the pointwise Jacobi smoother error estimator $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$.
Figure 5.5: Convergence of $\mathcal{P}_{p}$-AFEM driven by Jacobi (Left) and Gauss–Seidel error estimators (Right) for the Helmholtz Equation with $k=40\pi$ and estimator/error shown in the parenthesis.
Figure 5.6: Effectivity ratios $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}/\|u-u_{h}\|_{1,k}$ (top) and $\|S^{m}_{\mathcal{P}_{p}}r\|_{1,k}/\|u-u_{h}\|_{1,k}$ (bottom) for the Helmholtz equation.
@@ 6 Concluding Remarks
This paper develops a practical framework connecting a posteriori error estimates with smoothers of iterative linear solvers. By leveraging smoothers of available preconditioners and linear solvers, we have derived a posteriori error estimates for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. For symmetric problems, reliability and efficiency of smoother-type error estimators have been theoretically verified with the help of two-level stable decompositions and saturation assumptions. In addition, extensive numerical experiments demonstrate their robustness with respect to the polynomial degree.
@@ Acknowledgments
Y. L. would like to thank Prof. Shihua Gong for helpful discussion about iterative solvers for the Helmholtz equation. The authors would like to thank the anonymous referees for constructive remarks that improved the quality and presentation of this paper.